# ADT4Coupons: An Innovative Framework for Sequential Coupon Distribution in E-commerce

### v1

@@ Abstract
Coupon distribution is a critical marketing strategy used by online platforms to boost revenue and enhance user engagement. Regrettably, existing coupon distribution strategies fall far short of effectively leveraging the complex sequential interactions between platforms and users. This critical oversight, despite the abundance of e-commerce log data, has precipitated a performance plateau. In this paper, we focus on the scene that the platforms make sequential coupon distribution decision multiple times for various users, with each user interacting with the platform repeatedly. Based on this marketing scenario, we propose a novel marketing framework, named Aligned Decision Transformer for Coupons (ADT4Coupons), to directly devise coupon distribution policy for long-term revenue boosting. ADT4Coupons enables optimized online decision-making in a variety of real-world marketing scenarios. It achieves this by seamlessly integrating three key characteristics, general scenarios, sequential modeling with more comprehensive historical data, and efficient iterative updates within a unified framework. Furthermore, empirical results on real-world industrial dataset, alongside public and synthetic datasets demonstrate the superiority of our framework.
Keywords Sequential Decision-making  $\cdot$ Coupon Distribution  $\cdot$ Marketing Campaign
1 Introduction
In the highly competitive online ecosystem, coupon distribution has emerged as a central marketing strategy for online platforms looking to amplify user engagement and optimize revenue streams. Coupons take many forms (as shown in Figure 2), from traditional discounts to service upgrades and cash rewards, all designed to incentivize consumption across diverse sectors. For instance, Booking’s promotional activities can enhance user satisfaction Goldenberg et al. (2020); Albert and Goldenberg (2022); Airbnb employs dynamic pricing Ye et al. (2018); Uber’s promotions encourage users to adopt new products Du et al. (2019); Taobao’s discounts increase user activity Zhao et al. (2019); Zhang et al. (2021); Kuaishou offers cash rewards to stimulate user retention Ai et al. (2022); and Meituan utilizes spend&save discounts to incentivize purchases Zhou et al. (2023); Huang et al. (2024); Zhou et al. (2024).
Figure 1: Complex and Meaningful Coupon Distribution. Coupons take many forms, from traditional discounts to service upgrades and cash rewards, all designed to incentivize consumption across diverse sectors.
Figure 2: Single-Round Scenarios. Top: RL methods model the interactions with single user as Markov process. Below the timeline: Two-stage and DFL methods decide on coupon distribution for single users, based on current state alone.
@@ 1 Introduction
While prior research has explored counterfactual policy optimization from logged bandit feedback Swaminathan and Joachims (2015a, b); Joachims et al. (2018), the bandit approach relies on online exploration, rendering it unsuitable for coupon distribution in e-commerce scenarios where decisions, once made, are difficult to reverse. In contrast, more recent studies acknowledge the importance of pre-training models with logged data before deployment, representing a significant improvement. However, they typically focus on scenarios where the platform makes a single-round coupon allocation decision for each user interaction according to the current state. The prevailing two-stage approach dominates coupon allocation research Hansotia and Rukstales (2002); Zhao et al. (2019); Goldenberg et al. (2020), comprising: (1) machine learning (ML) prediction of user response to different coupons, and (2) optimization research (OR) for revenue maximization. However, decoupling the process into two stages leads to error accumulation, and the two-stage goals cannot be fully aligned. Recent end-to-end frameworks, including decision-focused learning (DFL) Zhou et al. (2023, 2024) and reinforcement learning (RL) Xiao et al. (2019); Zhang et al. (2021), attempt to address this limitation. However, these approaches face inherent performance limitations when constrained to this single-round interaction paradigm, as illustrated in Figure 2. In these settings, the potential for leveraging the richer sequential dependencies in user behavior is significantly curtailed, leading to a performance plateau.
Unlike these single-round paradigms, our work, while inheriting the benefits of an end-to-end framework, embraces the sequential nature of user engagement within e-commerce platforms. We propose an efficient framework, named Aligned Decision Transformer for Coupons (ADT4Coupons) to directly devise coupon distribution policy for long-term revenue boosting. Critically, while demonstrated here for coupon distribution, ADT4Coupons’s general framework empowers exploration of constrained optimization scenarios across many applications. Our contributions are threefold:
@@
Efficient Framework. ADT4Coupons framework offers two key advantages: (1) by explicitly modeling the series of user interactions and utilizing historical e-commerce data, ADT4Coupons captures valuable temporal dynamics overlooked by single-round approaches; and (2) by reformulating sequential decision-making as an auto-regressive prediction task, we leverage causal-Transformer structure to effectively capturing both immediate and multi-round causal relationships.Meanwhile, this structure reduces the framework’s reliance on the statistical premise of unbiased data, which serves as the training set foundation for two-stage methods and DFL based on uplift models. This allows for efficient, timely, and cost-effective iterative updates, contributing to sustained optimal performance in dynamic online environments.
@@
2 Further Related Work
Coupon distribution requires estimating user responses to different coupon strategies, a task known as counterfactual evaluation. A key metric is the conditional average treatment effect (CATE), which quantifies the impact of a specific coupon on outcomes like conversion rates, considering individual user characteristics Devriendt et al. (2018). Uplift modeling, a field dedicated to CATE estimation, has produced various machine learning techniques, often trained on randomized controlled trial (RCT) data Athey and Imbens (2015). The trained models then predict individual-level causal effects for arbitrary interventions. Notable CATE estimation techniques include meta-learners Hansotia and Rukstales (2002); Athey and Imbens (2015); Künzel et al. (2019); Nie and Wager (2021), ranking-based methods Kuusisto et al. (2014); Betlei et al. (2021), representation learning Johansson et al. (2016); Shi et al. (2019); Yao et al. (2018), and causal forests Ai et al. (2022); ATHEY et al. (2019); Wager and Athey (2018); Zhao et al. (2017). However, existing CATE formulations primarily address single-treatment effects. In contrast, our problem requires modeling temporal treatment dependencies, where user responses depend not only on the current coupon but also on historical coupon exposures. This introduces complex sequential counterfactual estimation challenges that traditional uplift models cannot address.
To tackle this challenge, we leverage the Transformer architecture Vaswani (2017), which has demonstrated remarkable capabilities in sequential modeling across diverse domains including natural language processing Keskar et al. (2019); Dai et al. (2019), computer vision Parvaiz et al. (2023), and reinforcement learning (RL) Chen et al. (2021); Zheng et al. (2022); Wu et al. (2023). While prior RL applications focus on unconstrained optimization of cumulative rewards, we adapt the Decision Transformer (DT) framework Chen et al. (2021) for constrained optimization problems. Our modification enables handling of budget constraints while preserving DT’s ability to model sequential dependencies in decision-making.
3 Coupon Distribution Problem
3.1 Model on Coupon Distribution
@@
An e-commerce platform has $N$ users. Given the requirement that all users register on the platform before purchasing, $N$ is a known parameter. Over a finite time horizon, users interact with the platform for up to $T$ times, referred to as $T$ rounds. Upon receiving a user interaction request, the platform must make a real-time decision regarding coupon issuance. Critically, once a coupon is granted, it cannot be revoked. Over the $T$ rounds, the platform has a marketing budget $B$, which is the total amount it is willing to spend on issuing coupons to all users. The platform offers $K$ types of coupons, such as $\{2,5,10,15,18\}$ dollar coupons for $K=5$, and so forth. Let $[N],[T],[K]$ denote the sets of users, rounds, and coupons, respectively.
@@
3.2 Dual Problem Formulation
Let $\boldsymbol{x}$ represent the coupon distribution strategy, $R(\boldsymbol{x})$ denote the total revenue obtained by the platform under this strategy, and $C(\boldsymbol{x})$ represent the total amount of coupons subsidized by the platform under this strategy. To obtain the optimal distribution strategy within the budget constraint, we first introduce $\lambda$ as the dual variable for the constraint $\sum_{i\in[N],t\in[T],k\in[K]}{c_{itk}x_{itk}}\leq B$. Next, we define the Lagrangian function $L(\boldsymbol{x},\lambda)$ as
$L(\boldsymbol{x},\lambda)=\sum_{i\in[N],t\in[T],k\in[K]}{r_{itk}x_{itk}}-\lambda\left(\sum_{i\in[N],t\in[T],k\in[K]}{c_{itk}x_{itk}}-B\right)$
The objective function of the dual problem is the maximum value of the Lagrangian function with respect to $\boldsymbol{x}$, i.e.,
$\min_{\lambda\geq 0}{\max_{0\leq x_{itk}\leq 1}{L(\boldsymbol{x},\lambda)}}=\min_{\lambda\geq 0}{\max_{0\leq x_{itk}\leq 1}{\left(R(\boldsymbol{x})-\lambda\left(C(\boldsymbol{x})-B\right)\right)}}$ (4)
@@ 3.2 Dual Problem Formulation
However, the platform can only observe the actual revenue/cost generated after a specific coupon decision and the counterfactual data is inherently unattainable. Furthermore, user responses are sequential. Previous coupon distribution decisions influence subsequent users’ coupon responses, making counterfactual prediction challenging. Therefore, the online linear programming problem cannot be directly solved.
4 Our Framework
To address the above challenge, we propose our framework named Aligned Decision Transformer for Coupons (ADT4Coupons), a direct policy learning approach involving training a model to determine the optimal sequential policy $\boldsymbol{x_{i}}$ for each user $i$ given $\lambda$, then leveraging a $\lambda$ optimization algorithm and model-parallel inference to derive the optimal overall strategy $\boldsymbol{x}$.
4.1 Trajectory Representation
@@
Given a dataset in the format (user-id, user-features as state, treatment as action, cost, reward, time) and additional $\lambda$ randomized in the interval $(0,1)$, we aggregate the dataset by user-id and then sort the data with each user’s group in chronological order. We define $RTG_{i_{t}}=\sum_{t^{\prime}=t}^{T}\gamma r_{it}$ Chen et al. (2021) and $CTG_{i_{t}}=\sum_{t^{\prime}=t}^{T}c_{it}$. The training dataset is pre-processed into trajectory data of size $N$ :
$\tau_{i\in[N]}=[(s_{i_{0}},a_{i_{0}},CTG_{i_{0}},RTG_{i_{0}},\lambda_{i}),\ldots,[(s_{i_{t}},a_{i_{t}},CTG_{i_{t}},RTG_{i_{t}},\lambda_{i}),\ldots,(s_{i_{T}},a_{i_{T}},CTG_{i_{T}},RTG_{i_{T}},\lambda_{i})]$
As mentioned above, the dual variable $\lambda$ reflects the constraint on our marketing budget $B$. To improve the model’s ability to generalize across diverse budget levels, we randomize $\lambda$ ten times for each trajectory sample. This effectively expands our training dataset tenfold, promoting more robust learning.
@@
4.2 Aligned Decision Transformer
Having established the training dataset, we adapt the Decision Transformer model and align it for direct coupon distribution policy learning, as summarized in Figure 3 and Algorithm 1.
Our model leverages two key advantages of the casual transformer structure, the ability to handle sequential data effectively via its transformer architecture and the capability to evaluate counterfactual outcomes through its casual structure.
To integrate constraint information, we augment the input space with the parameters ctgctg, representing constraint types. Furthermore, we enhance the original time embedding $Embed(t)$ by incorporating the dual variable embedding $Embed(\lambda)$, as shown in line 6-9 of Algorithm 1. This enhancement effectively encoded constraint information alongside temporal signals for each parameter. Subsequently, these augmented inputs are processed through the Casual Transformer blocks, with the internal details visualized in the left panel of Figure 3.
@@ 4.2 Aligned Decision Transformer
4.3 Overall Strategy
Algorithm 2 Overall Algorithm of ADT4Coupons
@@ 4.3 Overall Strategy
5:     $\text{penalty}\leftarrow\lambda\cdot\max(C(\boldsymbol{x})-B,0)$
@@
Based on the trained model to determine the optimal sequential policy $\boldsymbol{x_{i}}$ for each user $i$ given $\lambda$, we combine it with the $\lambda$ optimization algorithm, obtaining the ADT4Coupons framework to derive the optimal overall strategy $\boldsymbol{x}$, as summarized in Algorithm 2.
@@
Unlike prior single-round approaches, ADT4Coupons, as a sequential modeling framework, unlocks the potential to extract complex user-platform interaction patterns from rich datasets. Critically, while demonstrated here for coupon distribution, our general framework empowers exploration of constrained optimization scenarios across a spectrum of applications. Furthermore, its significant scalability allows for straightforward integration of external components, as detailed in the supplementary material.
5 Experiment
[…]
5.1 Experimental Settings
5.1.1 Dataset
CRITEO-UPLIFT v2 (CU). The dataset is provided by Criteo AI Labs at the AdKDD’18 workshop Diemert et al. (2018) and comprises 13.9 million samples collected from Randomized Controlled Trial (RCT). Each sample contains 12 features, a binary treatment indicator, and two response labels, namely visit and conversion. We follow the approach in Zhou et al. (2023), extracting user feature columns ‘f0-f11’ as states, ‘treatment’ as actions, ‘visit’ label as costs and ‘conversion’ label as generated revenue.
MT-LIFT (MT). Collected from two months of intelligent coupon marketing scenarios for food delivery in the Meituan appHuang et al. (2024), the dataset contains nearly 5.5 million samples, which include 99 features, treatment space with 5 options, and also two response labels: click and conversion. Similar to the processing method of the CU dataset, we process user feature columns ‘f0-f98’ as states, ‘treatment’ as actions, ‘click’ label as costs and ‘conversion’ label as generated revenue.
Synthetic Dataset (SY). We have developed a synthetic dataset generation pipeline inspired by Yan et al. (2023). The dataset consists of 10,000 simulated users. Each user interaction spans 10 timesteps, with an action space of four treatments applied randomly with equal probabilities at each timestep. The user responds based on the response probability, and corresponding cost and reward is generated when the user responds to the coupon. Gaussian noise is used to simulate state transitions.
ByteDance Marketing Data (BD). To realistically evaluate the performance of our model in real-world industrial scenarios, we incorporate a recent dataset originating from ByteDance, a global leading technology company with a diverse range of popular platforms. User features included in this dataset are anonymized to ensure data privacy and compliance. The real-world marketing dataset consists of millions of samples, including anonymized user features as states, treatments as actions, ‘cost’ label as costs and ‘is paid’ label as revenue.
For all datasets, we randomly split them separately, using 80% of the samples for training and 20% for testing. After initial data cleaning to remove anomalies, we aggregate the data by user ID. In cases where user ID was missing, we select a relatively stable user feature as the aggregation key. Within each aggregated group, we subsequently sort the data chronologically to generate trajectory samples.
5.1.2 Evaluation Metrics
We employ three metrics to evaluate the effectiveness of our experiments.
  • Return on Investment (ROI) ($\frac{R(\boldsymbol{x})-C(\boldsymbol{x})}{C(\boldsymbol{x})}$): The emphasis on the ROI indicator is paramount. Without proper consideration of cost, the strategy risks favoring marketing activities targeting users with both high returns and high associated expenditures. It leads to a suboptimal outcome and inefficient resource allocation.
5.2 Evaluation Experiments
5.2.1 Baseline Algorithms
  • Casual ForestAthey and Wager (2021): This baseline model, commonly used in marketing scenarios, is trained on historical data to predict individual treatment effects (ITE). At each decision point, the model estimates the treatment effect of offering a coupon and allocates it only if the predicted effect is positive.
  • Bandit Chapelle and Li (2011): We employ a multi-arm bandit algorithm based on Thompson Sampling (TS). This algorithm leverages Bayesian methods to construct a belief over the reward distribution of actions.
  • Constrained Markov Decision Process (CMDP) Xiao et al. (2019): For different $\lambda$ values, the CMDP model needs to be retrained. Hence, we established three CMDP baselines, CMDP25, CMDP50 and CMDP75, by retraining the core policy network with $\lambda$ values of 0.25, 0.5, and 0.75, respectively.
  • Decision Focused Causal Learning (DFCL) Zhou et al. (2024): The DCFL baseline is a state-of-the-art method that claims to achieve breakthrough performance by building upon two-stage approaches. It is an end-to-end framework, aiming to align the prediction objective with the ultimate decision-making goal.
@@ 5.2.1 Baseline Algorithms
5.2.2 Experimental Results
To establish a clear baseline, we select the Forest algorithm as our benchmark model. Figure 4 presents the revenue difference between evaluated models and the casual forest benchmark. We choose Forest due to its widespread adoption in marketing platforms, known for its robust causal learning capabilities and the generation of effective coupon distribution strategy. It is an appropriate and meaningful point of comparison.
Table 1: Performance Comparison Across Datasets and Baselines
Dataset Metric Ours DFCL Bandit CMDP25 CMDP50 CMDP75 Forest
@@ 5.2.2 Experimental Results
The experimental results in Figure 4 demonstrate that our proposed ADT4Coupons framework gains higher revenue compared to baselines across all datasets and budget levels, with two exception: the medium budget level in the SY dataset and MT dataset. ADT4Coupons achieves an average revenue improvement of 3.60%. Furthermore, Table 1 highlights ADT4Coupons’s superior performance in terms of both average ROI and BARate across the tested budget range. This overall superior performance emphasizes our framework’s strong strategy learning capabilities, robust causal reasoning, and adaptability to diverse dataset characteristics.
@@
5.3 Ablation Study
Figure 5: Ablation Study on MT Dataset.
5.3.1 Ablation Design
@@
  • No-$\widehat{RTG}$: We utilize the original Transformer model Vaswani (2017) without the $\widehat{RTG}$ parameter to evaluate its impact on the model’s capacity to pursue long-term revenue, as opposed to immediate rewards.
5.3.2 Experimental Results
@@
Comparing our framework with No-constraint model reveals that removing constraint-related variables leads to a diminished awareness of budgetary limitations. The model fails to adequately respond to budgetary changes. Comparing ADT4Coupons to model without $RTG$ shows that, when the budget is not sufficiently generous, the absence of $RTG$ causes the model to marginally exceed the budget in pursuit of immediate revenue gains. Furthermore, when budget is ample, the absence of RTG guidance restricts overall revenue generation, rendering its performance inferior to ADT4Coupons.
5.4 Inference Speed Analysis
Motivated by the demands of real-world applications, we conduct inference efficiency experiments to validate the runtime performance of our framework. The results are presented below.
Table 2: Model Inference Time

### v2

# SACO: Sequence-Aware Constrained Optimization Framework for Coupon Distribution in E-commerce

@@ Abstract
Coupon distribution is a critical marketing strategy used by online platforms to boost revenue and enhance user engagement. Regrettably, existing coupon distribution strategies fall far short of effectively leveraging the complex sequential interactions between platforms and users. This critical oversight, despite the abundance of e-commerce log data, has precipitated a performance plateau. In this paper, we focus on the scene that the platforms make sequential coupon distribution decision multiple times for various users, with each user interacting with the platform repeatedly. Based on this scenario, we propose a novel marketing framework, named Sequence-Aware Constrained Optimization (SACO) framework, to directly devise coupon distribution policy for long-term revenue boosting. SACO framework enables optimized online decision-making in a variety of real-world marketing scenarios. It achieves this by seamlessly integrating three key characteristics, general scenarios, sequential modeling with more comprehensive historical data, and efficient iterative updates within a unified framework. Furthermore, empirical results on real-world industrial dataset, alongside public and synthetic datasets demonstrate the superiority of our framework.
Introduction
In the highly competitive online ecosystem, coupon distribution has emerged as a central marketing strategy for online platforms looking to amplify user engagement and optimize revenue streams. Coupons take many forms (as shown in Figure 1), from traditional discounts to service upgrades and cash rewards, all designed to incentivize consumption across diverse sectors. For instance, Booking’s promotional activities can enhance user satisfaction (Goldenberg et al. 2020; Albert and Goldenberg 2022); Airbnb employs dynamic pricing (Ye et al. 2018); Uber’s promotions encourage users to adopt new products (Du et al. 2019); Taobao’s discounts increase user activity (Zhao et al. 2019; Zhang et al. 2021); Kuaishou offers cash rewards to stimulate user retention (Ai et al. 2022); and Meituan utilizes spend&save discounts to incentivize purchases (Zhou et al. 2023; Huang et al. 2024; Zhou et al. 2024).
Figure 1: Complex and Meaningful Coupon Distribution.
@@ Introduction
While prior research has explored counterfactual policy optimization from logged bandit feedback (Swaminathan and Joachims 2015b, a; Joachims et al. 2018), the bandit approach relies on online exploration, rendering it unsuitable for coupon distribution in e-commerce scenarios where decisions, once made, are difficult to reverse. In contrast, more recent studies acknowledge the importance of pre-training models with logged data before deployment, representing a significant improvement. However, they typically focus on scenarios where the platform makes a single-round coupon allocation decision for each user interaction according to the current state. The prevailing two-stage approach dominates coupon allocation research (Hansotia and Rukstales 2002; Zhao et al. 2019; Goldenberg et al. 2020), comprising: (1) machine learning (ML) prediction of user response to different coupons, and (2) optimization research (OR) for revenue maximization. However, decoupling the process into two stages leads to error accumulation, and the two-stage goals cannot be fully aligned. Recent end-to-end frameworks, including decision-focused learning (DFL) (Zhou et al. 2023, 2024) and reinforcement learning (RL) (Xiao et al. 2019; Zhang et al. 2021), attempt to address this limitation. However, these approaches face inherent performance limitations when constrained to this single-round interaction paradigm, as illustrated in Figure 2. The top part visualizes the RL methods modeling the interactions with single user as Markov processes. In contrast, the bottom part shows Two-stage and DFL methods making decisions based on current state alone. In these settings, the potential for leveraging the richer sequential dependencies in user behavior is significantly curtailed, leading to a performance plateau.
Figure 2: Single-Round Scenarios in Existing Research.
Unlike these single-round paradigms, our work, while inheriting the benefits of an end-to-end framework, embraces the sequential nature of user engagement within e-commerce platforms. We propose an efficient framework named Sequence-Aware Constrained Optimization (SACO) framework to directly devise coupon distribution policy for long-term revenue boosting. Critically, while demonstrated here for coupon distribution, SACO’s general framework empowers exploration of constrained optimization scenarios across many applications. Our contributions are threefold:
@@
Efficient Framework. SACO framework offers two key advantages: (1) by explicitly modeling the series of user interactions and utilizing historical e-commerce data, SACO captures valuable temporal dynamics overlooked by single-round approaches; and (2) by reformulating sequential decision-making as an auto-regressive prediction task, we leverage causal-Transformer structure to effectively capturing multi-round causal relationships. Meanwhile, this structure reduces the framework’s reliance on the statistical premise of unbiased data, which serves as the training set foundation for two-stage methods and DFL based on uplift models. This allows for efficient, timely, and cost-effective iterative updates, contributing to sustained optimal performance in dynamic online environments.
@@
Further Related Work
Coupon distribution requires estimating user responses to different coupon strategies, a task known as counterfactual evaluation. A key metric is the conditional average treatment effect (CATE), which quantifies the impact of a specific coupon on outcomes like conversion rates, considering individual user characteristics (Devriendt et al. 2018). Uplift modeling, a field dedicated to CATE estimation, has produced various machine learning techniques, often trained on randomized controlled trial (RCT) data (Athey and Imbens 2015). The trained models then predict individual-level causal effects for arbitrary interventions. Notable CATE estimation techniques include meta-learners (Hansotia and Rukstales 2002; Athey and Imbens 2015; Künzel et al. 2019; Nie and Wager 2021), ranking-based methods (Kuusisto et al. 2014; Betlei et al. 2021), representation learning (Johansson et al. 2016; Shi et al. 2019; Yao et al. 2018), and causal forests (Ai et al. 2022; ATHEY et al. 2019; Wager and Athey 2018; Zhao et al. 2017). However, existing CATE formulations primarily address single-treatment effects. In contrast, our problem requires modeling temporal treatment dependencies, where user responses depend not only on the current coupon but also on historical coupon exposures. This introduces complex sequential counterfactual estimation challenges that traditional uplift models cannot address.
To tackle this challenge, we leverage the Transformer architecture (Vaswani 2017), which has demonstrated remarkable capabilities in sequential modeling across diverse domains including natural language processing (Keskar et al. 2019; Dai et al. 2019), computer vision (Parvaiz et al. 2023), and reinforcement learning (RL) (Chen et al. 2021; Zheng et al. 2022; Wu et al. 2023). While prior RL applications focus on unconstrained optimization of cumulative rewards, we adapt the Decision Transformer (DT) framework (Chen et al. 2021) for constrained optimization problems. Our modification enables handling of budget constraints while preserving DT’s ability to model sequential dependencies in decision-making.
Coupon Distribution Problem
Model on Coupon Distribution
@@
An e-commerce platform has $N$ users. Given the requirement that all users register on the platform before purchasing, $N$ is a known parameter. Over a finite time horizon, users interact with the platform for up to $T$ times, referred to as $T$ rounds. Upon receiving a user interaction request, the platform must make a real-time decision regarding coupon issuance. Critically, once a coupon is granted, it cannot be revoked. Over the $T$ rounds, the platform has a marketing budget $B$, which is the total amount it is willing to spend on issuing coupons to all users. The platform offers $K$ types of coupons, such as $\{2,5,10\}$ dollar coupons for $K=3$, and so forth. Let $[N],[T],[K]$ denote the sets of users, rounds, and coupons, respectively.
@@
Dual Problem Formulation
Let $\boldsymbol{x}$ represent the coupon distribution strategy, $R(\boldsymbol{x})$ denote the total revenue obtained by the platform under this strategy, and $C(\boldsymbol{x})$ represent the total amount of coupons subsidized by the platform under this strategy. To obtain the optimal distribution strategy within the budget constraint, we first introduce $\lambda$ as the dual variable for the constraint $\sum_{i\in[N],t\in[T],k\in[K]}{c_{itk}x_{itk}}\leq B$. Next, we define the Lagrangian function $L(\boldsymbol{x},\lambda)$ and the objective function of the dual problem is the maximum value of the Lagrangian function with respect to $\boldsymbol{x}$, i.e.,
$\min_{\lambda\geq 0}{\max_{0\leq x_{itk}\leq 1}{L(\boldsymbol{x},\lambda)}}$ (4)
$=$ $\min_{\lambda\geq 0}{\max_{0\leq x_{itk}\leq 1}{\left(R(\boldsymbol{x})-\lambda\left(C(\boldsymbol{x})-B\right)\right)}}$ (5)
@@
Assuming all responses are known, the problem simplifies to a classical multiple choice knapsack problem, yielding an approximation ratio near 1(Zhou et al. 2023, 2024). However, the platform can only observe the actual revenue or cost after a specific coupon decision and the counterfactual data is inherently unattainable. Furthermore, user responses are sequential. Previous coupon distribution decisions influence subsequent users’ coupon responses, making counterfactual prediction challenging. Therefore, the online linear programming problem cannot be directly solved.
Our Framework
To address the above challenge, we propose our framework named Sequence-Aware Constrained Optimization (SACO) framework, a direct policy learning approach involving training a model to determine the optimal sequential policy $\boldsymbol{x_{i}}$ for each user $i$ given $\lambda$, then leveraging a $\lambda$ optimization algorithm and model-parallel inference to derive the optimal overall strategy $\boldsymbol{x}$.
Trajectory Representation
@@
Given a dataset in the format (user-id, user-features as state, treatment as action, cost, reward, time) and additional $\lambda$ randomized in the interval $(0,1)$, we aggregate the dataset by user-id and then sort the data with each user’s group in chronological order. We define $RTG_{i_{t}}=\sum_{t^{\prime}=t}^{T}\gamma r_{it}$ (Chen et al. 2021) and $CTG_{i_{t}}=\sum_{t^{\prime}=t}^{T}c_{it}$. The training dataset is pre-processed into trajectory data of size $N$ :
$\tau_{i\in[N]}=$ $[(s_{i_{0}},a_{i_{0}},CTG_{i_{0}},RTG_{i_{0}},\lambda_{i}),\ldots,$
$[(s_{i_{t}},a_{i_{t}},CTG_{i_{t}},RTG_{i_{t}},\lambda_{i}),\ldots,$
$(s_{i_{T}},a_{i_{T}},CTG_{i_{T}},RTG_{i_{T}},\lambda_{i})]$
As mentioned above, the dual variable $\lambda$ reflects the constraint on the budget $B$. To improve the model’s ability to generalize across diverse budget levels, we randomize $\lambda$ ten times for each trajectory sample. This expands our training dataset tenfold, promoting more robust learning.
@@
Constrained-Optimization Decision Transformer
Having established the training dataset, we adapt the Decision Transformer model and align it for direct policy learning, as summarized in Figure 3 and Algorithm 1.
Our model leverages two key advantages of the causal transformer structure, the ability to handle sequential data effectively via its transformer architecture and the capability to evaluate counterfactual outcomes through its causal structure(Nichani et al. 2024).
To integrate constraint information, we augment the input space with the parameters ctgctg, representing constraint types. Furthermore, we enhance the original time embedding $Embed(t)$ by incorporating the dual variable embedding $Embed(\lambda)$, as shown in line 6-9 of Algorithm 1. This enhancement effectively encoded constraint information alongside temporal signals for each parameter. Subsequently, these augmented inputs are processed through the Causal Transformer blocks, with the internal details visualized in the left panel of Figure 3.
@@
Overall Strategy
Algorithm 2 Overall Algorithm of SACO
@@
[…]
@@
Based on the trained model to determine the optimal sequential policy $\boldsymbol{x_{i}}$ for each user $i$ given $\lambda$, we combine it with the $\lambda$ optimization algorithm, obtaining the SACO framework to derive the optimal overall strategy $\boldsymbol{x}$, as summarized in Algorithm 2.
@@
Unlike prior single-round approaches, SACO, as a sequential modeling framework, unlocks the potential to extract complex user-platform interaction patterns from rich datasets. Critically, while demonstrated here for coupon distribution, SACO’s general framework empowers exploration of constrained optimization scenarios across a spectrum of applications. Furthermore, its significant scalability allows for straightforward integration of external components.
Experiment
@@
Dataset Metric SACO DFCL Bandit CMDP25 CMDP50 CMDP75 Forest
@@
Table 1: Performance Comparison Across Datasets and Baselines
[…]
Figure 5: Ablation Study on MT Dataset.
Experimental Settings
Dataset
CRITEO-UPLIFT v2 (CU). The dataset is provided by Criteo AI Labs at the AdKDD’18 workshop (Diemert et al. 2018) and comprises 13.9 million samples collected from Randomized Controlled Trial (RCT). Each sample contains 12 features, a binary treatment indicator, and two response labels, namely visit and conversion. We follow the approach in (Zhou et al. 2023), extracting user feature columns ‘f0-f11’ as states, ‘treatment’ as actions, ‘visit’ label as costs and ‘conversion’ label as generated revenue.
MT-LIFT (MT). Collected from two months of coupon marketing scenarios for food delivery(Huang et al. 2024), the dataset contains nearly 5.5 million samples, which include 99 features, 5 types of treatment, and two response labels: click and conversion. Similar to the CU dataset, we process user feature columns ‘f0-f98’ as states, ‘treatment’ as actions, ‘click’ as costs and ‘conversion’ as revenue.
Synthetic Dataset (SY). We have developed a synthetic dataset generation pipeline inspired by (Yan et al. 2023). The dataset consists of 10,000 simulated users. Each user interaction spans 10 timesteps, with an action space of four treatments applied randomly with equal probabilities at each timestep. The cost and reward is generated when the user responds to the coupon with the response probability. Gaussian noise is used to simulate state transitions.
Business Dataset (BD). To realistically evaluate the performance of our model in real-world industrial scenarios, we incorporate a recent business dataset originating from ByteDance, a global leading technology company with a range of popular platforms. User features included are anonymized to ensure data privacy and compliance. The marketing dataset consists of millions of samples, including anonymized user features as states, treatments as actions, ‘cost’ label as costs and ‘is paid’ label as revenue.
For all datasets, we randomly split them separately, using 80% of the samples for training and 20% for testing. After initial data cleaning to remove anomalies, we aggregate the data by user ID or relatively stable user feature. Within each aggregated group, we subsequently sort the data chronologically to generate trajectory samples.
Evaluation Metrics
  • Return on Investment (ROI) ($\frac{R(\boldsymbol{x})}{C(\boldsymbol{x})}$): The emphasis on the ROI indicator is paramount. Without proper consideration of cost, the strategy risks favoring marketing activities targeting users with both high returns and high associated expenditures. It leads to a suboptimal outcome and inefficient resource allocation.
Evaluation Experiments
Baseline Algorithms
  • Causal Forest(Athey and Wager 2021): This baseline model, commonly used in marketing scenarios, is trained on historical data to predict individual treatment effects (ITE). At each decision point, the model estimates the treatment effect of offering a coupon and allocates it only if the predicted effect is positive.
  • Bandit (Chapelle and Li 2011): We employ a multi-arm bandit algorithm based on Thompson Sampling (TS). This algorithm leverages Bayesian methods to construct a belief over the reward distribution of actions.
  • Constrained Markov Decision Process (CMDP) (Xiao et al. 2019): For different $\lambda$ values, the CMDP model needs to be retrained. Hence, we established three CMDP baselines, CMDP25, CMDP50 and CMDP75, by retraining the core policy network with $\lambda$ values of 0.25, 0.5, and 0.75, respectively.
  • Decision Focused Causal Learning (DFCL) (Zhou et al. 2024): The DCFL baseline is a state-of-the-art method that claims to achieve breakthrough performance by building upon two-stage approaches. It is an end-to-end framework, aiming to align the prediction objective with the ultimate decision-making goal.
Experimental Results
To establish a clear baseline, we select the Forest algorithm as our benchmark model. Figure 4 presents the revenue difference between evaluated models and the causal forest benchmark. We choose Forest due to its widespread adoption in marketing platforms, known for its robust causal learning capabilities. It is an appropriate and meaningful point of comparison.
The experimental results in Figure 4 demonstrate that our proposed SACO framework gains higher revenue compared to baselines across all datasets and budget levels, with two exception: the medium budget level in the SY dataset and MT dataset. SACO achieves an average revenue improvement of 3.60%. Furthermore, Table 1 highlights SACO’s superior performance in terms of both average ROI and BARate across the tested budget range. This overall superior performance emphasizes SACO’s strong strategy learning capabilities, robust causal reasoning, and adaptability to diverse dataset characteristics.
@@
Ablation Study
Ablation Design
@@
  • No-$\widehat{RTG}$: We utilize the original Transformer model (Vaswani 2017) without the $\widehat{RTG}$ parameter to evaluate its impact on the model’s capacity to pursue long-term revenue, as opposed to immediate rewards.
Experimental Results
@@
Comparing our framework with No-constraint model reveals that removing constraint-related variables leads to a diminished awareness of budgetary limitations. The model fails to adequately respond to budgetary changes. Comparing SACO to model without $RTG$ shows that, when the budget is not sufficiently generous, the absence of $RTG$ causes the model to marginally exceed the budget in pursuit of immediate revenue gains. Furthermore, when budget is ample, the absence of RTG guidance restricts overall revenue generation, rendering its performance inferior to SACO.
Inference Speed Analysis
Motivated by the demands of real-world applications, we conduct inference efficiency experiments to validate the runtime performance of SACO. The results are presented below.